### v2

# Short-wavelength mesophases in the ground states of core-softened particles in two-dimensions

Abstract
We describe the formation of short-wavelength mesophases in a two-dimensional core-softened particle system. By proposing a series of specific ansatz for each relevant phase, we performed a variational analysis to obtain the ground-state phase diagram. Our results reveal a variety of cluster lattice phases with distinct cluster orientations, alongside traditional two-dimensional Bravais lattices such as square, triangular, oblique, and rectangular structures, as well as other non-Bravais arrangements including honeycomb and kagome phases. We characterize in detail the ground-state phase transitions and identify coexistence regions between competing phases, capturing both first-order and continuous transitions. In addition, we highlight the crucial role of the competing length scales introduced by the hard-core repulsion in shaping the rich landscape of mesophases, emphasizing the interplay between intra-cluster structure and inter-cluster organization. This study provides a systematic framework that could support future investigations of classical thermal melting behavior or quantum phase transitions in similar cluster-forming systems.
  • July 2025
1 Introduction
The spontaneous emergence of spatially modulated phases constitutes a central topic in the study of complex many-body systems, encompassing a wide range of physical contexts. These phases often arise from the presence of competing interactions, which introduce frustration into the system [1, 2]. In classical systems, such frustration can give rise to a variety of modulated phases [3, 4], including cluster phases [5, 6, 7] termed also “clump” [8, 9] or “bubble” phases [10]; inverted-cluster phases [11], also known as “inverse micelle” [3], “hole” [12] or “anticlump” phases [13]; and stripe phases [14, 15, 16]. Such phases have been observed experimentally in systems as diverse as block copolymers [17], magnetic thin films [18, 19], vortex matter [20] and colloidal systems [21, 22, 23]. In the last case, even more exotic phases like sigma [24], quasicrystaline [25, 26] and cluster-quasicrystaline [27] phases have been observed in simulations of model systems.
While such structures arise in classical systems, in quantum systems the interplay between competing interactions and quantum fluctuations gives rise to even more exotic phases, including supersolids, as is the case of dipolar Bose-Einstein condensates observed both experimentally [28, 29, 30] and theoretically [31, 32], as is the case of soft-core bosons that display quantum cluster crystals phases [33, 34, 35, 36], where superfluidity may coexist with non-trivial spatial order [37, 38]. In this context, quantum quasicrystaline phases can be observed at finite temperature [39], and also at the ground-state level, either by mean of an external potential [40] or by careful tuning the pairwise interactions [41, 42], giving rise to exotic quantum states of matter [43, 44].
Among the class of interactions that drive self-assembly, ultra-soft repulsive potentials, characterized by a bounded core and a negative minimum in their Fourier transform, are responsible for the formation of cluster-crystalline states with particles that might overlap [5, 45, 46]. Typically these potentials produce the stabilization of cluster phases with an integer $n$ occupation number and isostructural transitions as the density is increased [47]. Notably, these are first order phase transitions, that occur through the development of coexistence regions of phases with $n$ and $n+1$ occupancy numbers. Hence, at the beginning of the coexistence region, droplets of the $n+1$ cluster phase nucleates and, as the density increases, these droplets grow until the system presents a pure $n+1$ cluster phase and the coexistence region ends.
When a hard-core repulsion or additional sharp length scale is introduced at the short distances, the tendency to form overlapping clusters can be significantly inhibited, paving the way for the formation of new modulated states with properties between solid and liquid phases, for which are sometimes called mesophases [3]. At higher densities, such modulated phases includes long-wavelength stripes, clusters and holes [2, 13, 12, 48]. At lower densities, however, these phases are characterized by short-wavelength modulations [49, 50, 51, 52, 53], and includes phases such as square lattices [54], as well as stripes of one particle width, kagome and honeycomb phases [24, 55, 56, 57], among many others. In particular, the cluster phases with a small occupation number $n$ are often identified as dimers for $n=2$ [58], trimers ($n=3$), tetramers ($n=4$) [59], and so forth.
The mechanism for the pattern formation stems from the competition between the ultra-soft component, which favors clustering, and the hard-core component, which penalizes such configurations [48, 60]. This combination of soft and hard-core repulsions is sometimes referred as core-softened or core-corona potentials as well as hard-core/soft-shoulder (HCSS) potentials [61, 49, 3]. As a result, systems may settle into phases where the stabilized structure arises from the balance of competing interactions rather than simple clustering — a behavior observed in both classical and quantum contexts [62].
In situations when the effective hard-core size is much lower than the spacing of the lattice formed by the arrangement of the clusters themselves, the particles belonging to a cluster develop an internal structure that minimizes the hard-core repulsion energy - typically a BCC or triangular lattice in three or two dimensions, respectively. Meanwhile, in this case, the long-distance physical properties of the cluster lattice remains essentially unaltered in comparison to the corresponding soft-core particle system [48]. This kind of behavior naturally disappears as soon as the size of the clusters becomes comparable to the distance between them. In such regime the intra and inter cluster interactions are not well separated by an energy scale and, consequently, the behavior of the system becomes much richer, which can be explored already at the ground state level [63]. This scenario is exactly the one explored in Refs. [60] and [64], where the authors investigated the properties of a self-assembled crystalline dimer phase in a regime where the hard-core size of the particles is comparable to the distance between dimers forming the lattice. Interestingly, the authors shows that the classical ground state is given by a “nematic” state in which dimers are oriented in a non-trivial direction with respect to the dimers lattice.
Within this context, our work aims to present a systematic study of the self-assembling process in a 2D soft-core cluster forming model frustrated by a hard-core interaction. In order to do so, we consider a two dimensional GEM-4 model combined with an inverse power-law potential, as an effective hard-core interaction, and compute the exact classical ground state phase diagram, using as running parameters the density and the relative intensity between the GEM-4 interaction and the short-range power-law repulsion. Our results shows that the soft-core scenario, when the transition between clusters with increasing occupancy occurs as the density is increased, is significantly modified even when the size of the hard-core is relatively small. To be specific, the maximum cluster occupancy attained decreases rapidly with the size of the hard-core while a plethora of new single particle phases, not present in the phase diagram of the GEM-4 model, are developed. Finally, we would like to remark that, differently from previously studied scenarios in which the interparticle distance is much smaller than the distance between clusters, in our case the orientational order of the clusters is a non-trivial feature of the self assembling process. This makes the orientational order of the cluster states to vary along the phase diagram.
The paper is organized as follows, in Sect. 2 we introduce the model studied and the method employed for the determination of the classical ground state of the system. The Sect. 3 shows the results obtained for the phase diagram of the system, initially implementing the direct energy minimization for pure phases only and, in a second moment, including the possibility of coexistence between different phases. Moreover, a detailed description of the phases stabilized by the system is presented making evident the rich physical behavior of the system studied. Finally, Sect. 4 is devoted to present our final remarks and conclusions.
2 Model and Methodology
2.1 Core-softened repulsive potential model
As mentioned in the previous section, we consider a two-dimensional system of classical interacting particles with a core-softened repulsive pair potential of the form:
$V(r)=\exp\left(-r^{\alpha}\right)+\frac{C}{r^{\gamma}},$ (1)
with $\alpha=4$ and $\gamma=6$. The first term in (1) represents the soft part of the pairwise potential, known as the generalized exponential model (GEM-$\alpha$), which is a bounded potential allowing for the formation of cluster-crystal ground states, where clusters are arranged on a triangular lattice and each cluster is formed by overlapping particles, with the occupation number increasing with density for $\alpha>2$ [65, 45, 47]. This ultra-soft repulsion is combined with an inverse-power-law potential, the second term in (1), which effectively acts as a hard-core repulsion, preventing particles from overlapping and thereby enabling the emergence of a variety of fine-structured ground states, partially characterized in this work.
The parameter $C$ controls the relative strength of the hard-core repulsion with respect to the ultra-soft interaction, thereby regulating the different spatial scales on which these two interactions act. For very small $C$ (in the limit $C\to 0$), the hard-core repulsion acts on very short length scales, and the ground state of overlapping particle clusters is replaced by one in which the clusters acquire a small size, composed of particles arranged in a triangular lattice, with the distance between clusters being much larger than the distance between particles within each cluster [48]. In the opposite limit, when $C\gtrsim 0.16$, the ultra-soft repulsion is dominated by the hard-core potential and the ground-state becomes the familiar triangular lattice.
In this work, we are interested in the intermediate regime where both repulsive interactions act on comparable scales and promote complex mesophases. For example, we expect the formation of ground states with a regular lattice of clusters, where the distance between clusters is comparable to the distance between particles inside a cluster, so that the internal structure and orientation of each cluster affect the arrangement of the overall lattice. It is precisely in this regime that mesophases proliferate, giving rise to an intricate phase diagram.
2.2 From generic to specific ansatz
In order to identify candidates for possible short-wavelength ground-states we have performed molecular dynamics simulations. These were carried out in a NVT ensemble through an overdamped Langevin equation in dimensionless form, given by:
$\dot{\textbf{r}}_{i}=-\sum_{j}\nabla_{i}V($ $\left|\textbf{r}_{i}-\textbf{r}_{j}\right|)+\sqrt{2T}\,\xi_{i}(t),$ (2)
where $T$ is the temperature of the thermal bath and $\xi_{i}(t)$ is a white noise with $\left<\xi_{i}(t)\right>=0$ and $\left<\xi_{i}(t)\xi_{j}(t^{\prime})\right>=\delta_{ij}\delta(t-t^{\prime})$. These simulations were performed using 400 particles, whose equations of motion were integrated using a stochastic RK4 algorithm [66], with the addition of the parallel tempering technique [67, 68] in order to overcome metastable states. These exploratory results are not presented here and have not limited in any aspect our results, but were instrumental to initiate the construction of the ground-state ansatz for each phase.
The simulations suggested that, in the low-density region of interest $\rho<2.5$, corresponding to short-wavelength ground states, the cluster phases have occupation numbers between 2 and 4. To explore as many possible cluster ground states as feasible with energy minimization, we began by considering a generic ansatz, illustrated for cluster 2 in Figure 1. Rather than assuming a triangular lattice of clusters, we allowed a general angle $\theta$ between primitive translation vectors and variable lattice parameters along each primitive direction, described by aa along the x-axis and $b$ along the $\theta$ direction. The distance between particles within a cluster is denoted $d_{1}$, oriented along angle $\phi_{1}$. Additionally, we allowed for a sublattice displaced from the first lattice by a distance $b_{2}$ at angle $\theta_{2}$, with its own cluster size and orientation described by $d_{2}$ and $\phi_{2}$. Similar general ansatz were constructed for clusters with occupation numbers 3 and 4.
Figure 1: Schematic representation of the general ansatz employed to describe cluster 2 ground-states. The variational parameters considered are highlighted in the figure. Parameters $a$, $b$ and $\theta$ characterize the lattice of clusters whereas $d_{1}$ and $\phi_{1}$ and $d_{2}$ and $\phi_{2}$ describe the structure of the dimers on each sub-lattice. Finally, the parameters $b_{2}$ and $\theta_{2}$ give us the relative position of the second sub-lattice respect to first one.
As with the molecular dynamics stage, the aim of the numerical minimization based on these general ansatz was exploratory. The parameters $a$, $b$ and $\theta$ are not constrained, and in principle can represent any of the five two-dimensional Bravais lattices (oblique, rectangular, centered rectangular, triangular, and square). These degrees of freedom, for example, enable a hexagonal lattice to adapt to significant cluster size distortions. The same general ansatz can also describe non-Bravais lattices, such as those identified in previous works [69, 60], where cluster 2 phases exhibit dimers with non-uniform orientations alternating along one lattice direction.
We can evaluate the energy per particle of the system by summing the potential energy corresponding to all pairs of particles in our system:
$\frac{E}{N}=\frac{1}{2ns}\left(\sum_{l,m}^{\infty}\sum_{q,q^{\prime}}^{n}\sum_{r,r^{\prime}}^{s}V(|\textbf{R}_{lm,r}+\textbf{u}_{q,r}-\textbf{u}_{q^{\prime},r^{\prime}}|)\right),$ (3)
where $n$ is the cluster size and $s$ is the number of sub-lattices (in our case $s=2$). Moreover, the set $\{\textbf{R}_{lm,r}\}$ and $\{\textbf{u}_{q,r}\}$ represent the positions of the center of the clusters and the positions of the particles in the cluster respect to its center for each sub-lattice $r$, respectively. The indexes $l$ and $m$ serve to identify a specific lattice vector once we specify the corresponding lattice vector basis. As an example, to reproduce the configuration in Figure 1, we should set:
$\textbf{R}_{lm,1}$ $\equiv l\textbf{a}+m\textbf{b},$ (4)
$\textbf{R}_{lm,2}$ $\equiv l\textbf{a}+m\textbf{b}+\textbf{b}_{2},$
where the set $\{\textbf{a},\textbf{b}\}$ and $\textbf{b}_{2}$ represent the oblique vector basis of the lattice and relative position of the second sublattice respectively (see Fig. 1).
The lattice positions given by these vectors will be then decorated by the cluster structure, which can be described by a finite size triangular lattice. For the cluster 2 geometry, for instance, the positions can be parametrized as:
$\textbf{u}_{q,r}=\pm\frac{d_{r}}{2}\left\{\cos\phi_{r},\sin\phi_{r}\right\},\\$ (5)
where the $r$ index identifies the corresponding sub-lattice (1 or 2), whereas $d_{r}$ and $\phi_{r}$ represent the corresponding dimer size and orientation.
For the numerical minimization, we used the Sequential Quadratic Programming (SQP) solver for nonlinearly constrained problems from the ALGLIB library, version 3.16 [70]. The optional use of constraints helped avoid misidentification of phases. One issue with the general ansatz is that it cannot always sharply resolve phase boundaries or transitions, since it can adapt to different phases. For example, the dimer phase ansatz can also describe a triangular crystal. Another challenge is the large number of variational parameters, which requires testing different initial conditions to deal with the many metastable states. Nevertheless, combined with the molecular dynamics guidance, this approach allowed us to systematically identify the phases present in the system.
We identified traditional Bravais lattice phases with occupancy number $n=1$, such as triangular and square lattices, as well as single-particle rectangular and oblique lattices, corresponding in mesophase terminology to stripe phases with unit width. Their specific ansatz are shown in the first row of Figure 2.
Figure 2: Geometric representation of the different ansatz considered for each phase. The occupation number increases from the top to the bottom row. Each figure corresponds to a possible solution of our ground state phase diagram. The variational parameters, together with the main constraints on them, are presented for each configuration.
We also found two cluster 2 (dimer) phases, both in general assuming the form of oblique lattices. In the first, dimers are oriented along the line equidistant to neighboring dimers — described by the angle $\phi=\arctan(b\sin\theta/(a+b\cos\theta))$ — referred to here as the oblique cluster 2 lattice, or simply Cluster 2. In the second, the dimers align along a primitive direction of the oblique lattice, referred to as the oblique cluster 2-aligned lattice, or just Cluster 2a. Additionally, we identified a honeycomb lattice, corresponding to a cluster 2 “inverse micelle” or “hole”-type phase. The specific ansatz for these are illustrated in the second row of Fig. 2.
For cluster 3, we identified two phases: one with homogeneous trimer orientation, called the oblique cluster 3 lattice (or Cluster 3), and another with alternating trimer orientation along one primitive direction of the oblique lattice, designated the oblique cluster 3-alternated lattice (or Cluster 3a). Larger trimers can form a kagome lattice, corresponding to a cluster 3 hole-type phase. The third row of Fig.2 shows these, while the fourth row presents the ansatz for the oblique cluster 4 lattice (or Cluster 4 for simplicity), which is also capable to describe mesophases associated with cluster 4, such as stripes and holes, as the cluster distance decreases and clusters merge. For simplicity, we group all such cluster 4 mesophases under the general description of the oblique cluster 4 lattice phase.
This set of 11 different specific ansatz, represented in Figure 2, constitutes the basis for constructing the zero-temperature phase diagram of the system. Each was used to evaluate the energy as:
$\frac{E}{N}=\frac{1}{2n}\left(\sum_{l,m}^{\infty}\sum_{q,q^{\prime}}^{n}V(|\textbf{R}_{lm}+\textbf{u}_{q}-\textbf{u}_{q^{\prime}}|)\right),$ (6)
where $\textbf{R}_{lm}\equiv l\textbf{a}+m\textbf{b}$ and $\textbf{u}_{q+1}=\{d\cos(\phi+q\pi/3),d\sin(\phi+q\pi/3)\}$, for $q\geq 0$ and $\textbf{u}_{0}=\{0,0\}$. Here the index convention is the same as in Eq.(3). Note that in general, especially for larger clusters, it may be necessary to modify the $\textbf{u}_{q}$, however, up to cluster 4, our parametrization will accurately describe cluster geometries.
The only exception to the description given by eq. (6) is the Cluster 3a ansatz. In this case clusters alternate orientation along the b direction introducing a secondary sub-lattice in that direction. Nevertheless, in this case it is simple to realize that the total interaction energy of a trimer in both sub-lattices are equal. This results is straightforward if we observe that the trimers in both sub-latices are equivalent by a lattice reflection using a as symmetry axis. For this reason, to evaluate the total energy we simply consider the value obtained for a single type of trimer.
Furthermore, we should notice that at a given density $\rho$ the lattice parameters are not independent. In general, the density of particles can be evaluated as
$\rho={n_{c}}/{A},$ (7)
where $n_{c}$ is the number of particles inside the basic cell of the pattern having an area $A$. For example, for the oblique lattice the basic cell area is given as $A=ab\sin\theta$, which lead us to the conclusion:
$b=\frac{n_{c}}{a\rho\sin\theta}.$ (8)
Taking into account this kind of constraint, we have as independent variational parameters $a$, $\theta$ and $d$ for all cluster phases, $a$ and $\theta$ for the oblique phase and just $a$ for the rectangular phase. On the other hand, the triangular, square, honeycomb and kagome phases are fully determined by the density, without any minimization. In order to enforce each ansatz to represent a single type of configuration we should impose additional constraints to the variational parameters as presented in 2.
3 Results
3.1 Phase diagram
The systematic determination of the ground-state phase diagram was carried out over a wide range of densities $\rho\in[0.7,2.5]$ and values of the hard-core parameter $C\in[10^{-4},10^{-0.5}]$, expressed more naturally as $\ell_{C}\equiv-\log C$. Figure 3 summarizes the phase behavior of the system under these parameters. Several important observations emerge from this diagram.
0.51.01.52.02.53.03.54.00.91.21.51.82.12.4Refer to caption
$\ell_{C}$
$\rho$0.51.01.52.02.53.03.54.00.91.21.51.82.12.4Cluster 4Cluster 3aCluster 3Cluster 2aCluster 2KagomeHoneycombObliqueRectangularSquareTriangular
Figure 3: Ground state phase diagram for model presented in Eq. (1) using the density $\rho$ and relative strength of the hard-core interaction $\ell_{C}=-\log C$ as running parameters. The phase boundaries are represented with full lines in the case of first-order transitions, and by dashed lines in the case of continuous transitions.
To begin with, the overall structure of the phase diagram confirms that the addition of a hard-core repulsion to the GEM-4 ultra-soft interaction strongly enriches the phase behavior, suppressing the otherwise continuous sequence of cluster occupancies characteristic of the pure GEM-4 models [5, 45, 47]. Indeed, looking at the smallest values of $C$ explored (highest values of $\ell_{C}$), i.e., for weakest hard-core repulsion, one observes a sequence of cluster phases with occupancy numbers that grows with increasing density. This is expected from the known behavior of cluster crystal phases in systems with pure ultra-soft potentials, but with two remarkable differences. The first being the presence of more than one phase associated with a given occupation number. The second is that, as it will be discussed in detail shortly, the oblique lattice formed by the center of the clusters in some cases is closer to a triangular lattice than in others, depending on the occupancy number.
Naturally, for higher values of $C$ or lower values of $\ell_{C}$, the cluster size turns comparable to the lattice spacing between clusters, and clusters crystals phases give space for new single-particle phases, including square, rectangular, oblique, honeycomb and kagome phases. This demonstrates how the hard-core term disrupts particle overlap and favors unusual lattice structures in the context of ultra-soft pair potentials.
An interesting feature of the obtained phase diagram is the presence of both first-order (solid lines) and continuous (dashed lines) phase transitions. The first-order transitions generally appear between phases that are incommensurate with each other. These phases have distinct local symmetries or incompatible coordination motifs, preventing a continuous deformation from one to the other. As a consequence, the system exhibits abrupt changes in the particle distribution — typically involving changes in the local cluster occupancy or orientational order of the clusters — signaling a discontinuous jump in the absolute energy minimum configuration. A clear example is the transition from Cluster 2 to Cluster 3 phases, where the number of particles per cluster changes discretely.
By contrast, the dashed lines indicate continuous transitions, where the symmetry or local motif of one phase can be smoothly transformed into another by adjusting lattice parameters. For example, the Oblique lattice phase can continuously approach a Rectangular lattice phase as its lattice angle $\theta$ tends toward $\pi/2$, and the Rectangular lattice phase can evolve further into a Square lattice phase if its aspect ratio reaches unity. Similarly, a Cluster 2a phase can be continuously tuned into a Rectangular lattice phase as its internal distance $d$ grows. These pathways highlight the role of commensurability between phases, which defines whether a smooth transition is possible.
The presence of the Cluster 2a and Cluster 3a phases deserves emphasis. These phases appear as intermediates states between the more stable Cluster 2 and Cluster 3 phases. In particular, Cluster 2a exhibits a dimer structure oriented along one of the main directions of the lattice in a configuration that is reminiscent of a single particle width stripe phase, as is the case of the Oblique and Rectangular phases. In Cluster 3a, a more complex intra-cluster structure develop, suggesting a frustration between maximizing intra-cluster packing and optimizing the inter-cluster repulsion. Their relatively narrow stability windows reflect a delicate balance between these competing interactions, showing how the hard-core potential creates new local frustration absent in the purely ultra-soft case. In overall, the phase diagram illustrates how the introduction of a competing length scale through a hard-core term creates an expanded palette of mesophases in two dimensions, breaking the classical scenario of purely cluster-crystal states.
3.2 Ground-state configurations
In Figure 4, we show detailed representations of the ground-state configurations obtained by minimizing the energy within the specific ansatz defined previously. These snapshots illustrate the microscopic ordering of each phase and allow us to interpret their structural features in more depth.
[…]
Figure 4: Ground state configurations for selected values of $\rho$ and $\ell_{C}$ according to the phase diagram in Fig. 3.
For Cluster 2 and Cluster 2a [panels (a)–(c)], the configurations display a robust tendency to form dimer-like pairs of particles within each cluster. As $\ell_{C}$ decreases, the distance between particles inside each dimer increases, reflecting the growing influence of the hard-core repulsion. At the same time, the lattice parameter of the inter-cluster structure shrinks, leading to regimes where the intra- and inter-cluster distances become comparable. This is precisely the point where the orientational degree of freedom of the dimers becomes relevant. In the Cluster 2 phase, the dimers are oriented homogeneously along the direction of the midpoint between neighboring dimers, however such orientation varies continuously with $\rho$ and $\ell_{C}$, introducing a type of orientational order — or nematicity — that is analogous to behaviors seen in dimer-based molecular crystals and frustrated soft matter systems [69]. The possibility of tuning dimer orientations across the lattice marks a nontrivial extension of conventional cluster phases.
The Honeycomb and Kagome configurations [panels (d) and (h)] exhibit phases where the hard-core term dominates and prevents overlap, stabilizing open, high-density structures with high symmetry. These lattices are non-Bravais, but present well-defined local environments with uniform coordination numbers. Known for appearing in conditions of high frustration, such patterns can be interpreted as triangular lattices of holes, in a way that they minimize the repulsive energy, at the expense of bringing closer a number of particles. Finally, Cluster 3, Cluster 3a, and Cluster 4 phases [panels (i)–(l)] represent the higher cluster occupancy regimes explored. Their structures reveal triangular intra-cluster order, with the local motif approximately constant as $\rho$ grows. In Cluster 3a, the orientation of the clusters is not uniform across the lattice, alternating the direction of the trimers along one primitive direction. This highlights a frustration scenario where intra- and inter-cluster interactions are not easily satisfied.
40.045.050.055.060.02.02.53.03.54.0a)Refer to caption
$\theta$
$\ell_{C}$$\rho$(Cluster 2) 0.970.97(Cluster 2) 1.101.10(Cluster 2a) 1.421.42(Cluster 3a) 1.501.50(Cluster 3) $1.80$
54.057.060.063.00.91.21.51.82.12.4b)Refer to caption
$\theta$
[…]
Figure 5: Ground state angle value of the oblique lattice formed by the centroids of clusters. a) as a function of $\ell_{C}$ for different values of $\rho$, and b) as a function of $\rho$ for $\ell_{C}=3.99$. The triangular lattice case corresponds to $\theta=60^{\circ}$.
As discussed previously, cluster anisotropy distorts the lattice, even in the regime when the size of the clusters is relatively small in comparison with the lattice spacing of the cluster lattice. This behavior is however much stronger for more elongated clusters, such as when we have dimer phases. In Fig. 5, we explore the behavior of the equilibrium $\theta$ angle for different values of $\ell_{C}$ and $\rho$, characterizing how close to the triangular case ($\theta=60^{\circ}$) the general oblique cluster lattice is for the different cluster phases. As can be observed, for the Cluster 3 or 3a phases, where the internal symmetry of the clusters are compatible with the triangular lattice, the lattice angles are very close to $60^{\circ}$. On the other hand, for the Cluster 2 or 2a phases, the distortion is larger and converges very slowly to $60^{\circ}$ as $C\rightarrow 0$ ($\ell_{C}\to\infty$). This slow convergence indicate, as mentioned before, that even for quite small clusters the equilibrium cluster lattice is significantly anisotropic.
3.3 Coexistence regions
The phase diagram presented in Fig. (3) is not thermodynamically accurate as it does not take into account the possibility of phase coexistence as a mechanism to further minimize the energy of the system in the vicinities of discontinuous phase transitions. This aspect is simple to be corrected as the coexistence region associated to a first order phase transition can be calculated using the standard Maxwell construction. By matching the pressure $P$ and chemical potential $\mu$ of the competing phases at the beginning and end of the coexistence region, one guarantees that the first-order transition occurs through phase separation process rather than through an abrupt homogeneous configuration change.
The relevant thermodynamic quantities, pressure and chemical potential, can be calculated as
$P\equiv-\frac{\partial E}{\partial V}=\rho^{2}\frac{\partial\varepsilon}{\partial\rho},\quad\mu\equiv\frac{\partial E}{\partial N}=\varepsilon+\rho\frac{\partial\varepsilon}{\partial\rho},$ (9)
where $\varepsilon\equiv E/N$ stands for the energy per particle function, which is determined from the ground state energy calculation. Coexistence limits are then obtained by solving the following coupled equations
$\left\{\begin{aligned} P_{A}(\rho_{1})&=P_{B}(\rho_{2}),\\ \mu_{A}(\rho_{1})&=\mu_{B}(\rho_{2}),\end{aligned}\right$. (10)
where $\rho_{1}$ and $\rho_{2}$ marks the boundaries of the coexistence region along the density axis. The energy per particle through the coexistence region can then be evaluated by integrating the equation for the pressure, which yields
$\varepsilon_{\text{coex}}(\rho)=\varepsilon(\rho_{1})+P\left(\frac{1}{\rho_{1}}-\frac{1}{\rho}\right),$ (11)
allowing a consistent energetic comparison with the pure phases.
0.51.01.52.02.53.03.54.00.91.21.51.82.12.4Refer to caption
$\ell_{C}$
$\rho$0.51.01.52.02.53.03.54.00.91.21.51.82.12.4Cluster 4Cluster 3aCluster 3Cluster 2aCluster 2KagomeHoneycombObliqueRectangularSquareTriangularCoexistences
Figure 6: Ground-state phase diagram including coexistence between phases across first-order transitions. Grey regions represent the coexistence between neighboring phases along the density axis.
In Figure 6 we show the resulting phase diagram once we take into account the presence of coexistence regions. As can be observed this improvement modify significantly the thermodynamic phase diagram of the system, suppressing or eliminating entirely many of the pure phases in our system. This is the case, for instance, of the Cluster 2, Cluster 2a and Cluster 3a phases which presents significant coexistence regions at their phase boundaries. Moreover, we can verify the thermodynamic consistency of our calculations by examining the continuity of the extension of the coexistence regions when bifurcations in these regions occurs varying $\ell_{C}$. It is worth noticing that the extension of a coexistence region is an indicator of how strong a first order transition is, or in other words, how different the curves $\varepsilon(\rho)$ behave for each phase in the vicinity of the crossing density. Interestingly, the lower the values of $C$ the stronger are the first order transitions developed by the model. This indicates that the discontinuous transitions between configurations with different cluster occupancies are stronger than those in which single particle lattices reorganize in a different structure.
To complement the results for the phase diagram presented we show in Fig. 7 the behavior of the energy per particle, pressure and chemical potential increasing density at specific values of $\ell_{C}$. A closer look at this figure reveals the characteristic plateaus in the pressure and chemical potential expected in first order transitions. These plateaus are produced due to the presence of phase-separated domains, where the system locally minimizes its energy while maintaining mechanical and chemical equilibrium. Notably, the widths of most coexistence regions ($\Delta\rho$) are significant, suggesting that the transitions, while discontinuous, do not require extremely fine density control for experimental realization.
0.00.30.60.91.21.50.700.901.101.301.500.61.21.82.43.03.60.700.901.101.301.501.82.43.03.64.20.700.901.101.301.500.961.001.041.151.201.25Refer to caption
$\varepsilon$
$\rho$
$P$
[…]
$\mu$
$\rho$
0.00.51.01.52.02.50.71.11.51.92.30.02.04.06.08.00.71.11.51.92.31.02.03.04.05.06.00.71.11.51.92.3Refer to caption
$\varepsilon$
$\rho$
$P$
[…]
$\mu$
$\rho$
Figure 7: Energy per particle, pressure and chemical potential for different values of $\ell_{C}$ as a function of $\rho$. First and second row corresponds to $\ell_{C}=2.51$ and $\ell_{C}=3.99$, respectively.
Physically, the coexistence region tends to appear between phases with strongly mismatched local motifs, where continuous parameter deformations are energetically costly or forbidden. For example, coexistence is prominent between Cluster 2 and Cluster 3a, or between Cluster 3 and Cluster 4, whereas transitions involving Oblique and Rectangular structures have narrower coexistence intervals due to the possibility of smooth lattice deformations.
4 Conclusions
In this work, we have explored the ground-state phase diagram of a two-dimensional core-softened particle system, defined by the combination of the GEM-4 ultra-soft potential and an inverse power-law interaction with a decay exponent of 6 acting as a hard-core repulsion. At first, we conducted an exploration of the phase diagram using both molecular dynamics simulations and numerical energy minimization based on a general ansatz for $n$-occupied cluster phases with an independent orientational order in up to two sublattices. This strategy enabled us to identify the short-wavelength mesophases typically present in this model and to formulate a set of specific ansatz for each phase, respecting their symmetries while relying on the minimal set of parameters and constraints needed for their proper description.
By minimizing the energy of each phase with respect to its degrees of freedom, using a nonlinear constrained optimization technique, we have constructed the zero-temperature phase diagram using as running parameters the particle density and the relative strength of the hard-core potential compared to the ultra-soft repulsion. The resulting diagram reveals a wide variety of phases emerging in the intermediate regime where both interactions operate on comparable length scales. When the hard-core repulsion is much weaker, we observe cluster crystal phases with increasing occupancy as density grows. We identified that these cluster phases generally adopt oblique lattice arrangements, which tend to evolve toward triangular symmetry as the hard-core repulsion becomes negligible — in other words, as the cluster shape anisotropy becomes less significant due to its smaller size — although this convergence can be slow. A notable exception in this regard is the Cluster 3 phase, where the geometry of each cluster allows it to perfectly match the triangular lattice.
We have identified two distinct $n=2$ cluster (dimer) phases: the Cluster 2 phase, in which dimers align along a line equidistant from neighboring dimers, and the Cluster 2-aligned phase, where dimers are oriented along a primitive direction of the oblique lattice. We did not observe dimer phases with more complex orientational patterns, such as the alternated or “anti-nematic” dimer phases reported previously in similar two-dimensional core-softened systems with triangular-lattice constraints [69, 60]. However, we did observe such alternated behavior in the trimer case, the Cluster 3-alternated phase, where the compatibility between the internal cluster symmetry and the triangular lattice permits subtle energy optimization through variations in cluster orientation with minimal lattice distortion.
Beyond the cluster phases, as the hard-core repulsion becomes stronger, we have also identified single-particle phases, including triangular and square lattice arrangements, as well as phases associated with stripe-like mesophases (rectangular and oblique lattices) and hole-like mesophases (kagome and honeycomb lattices). The construction of the specific ansatz and the minimization of their energies not only allowed us to characterize and localize these phases in the phase diagram, but also to delineate coexistence regions near the phase transitions, thus mapping and classifying them as either discontinuous or continuous.
The detailed description of the short-wavelength region of the ground-state phase diagram presented in this work can establish the basis for future investigations of thermal and quantum fluctuation effects on mesophases. For example, this system could be an excellent candidate for studying the melting of anisotropic two-dimensional crystalline phases [71, 72], or, in the context of interacting bosons, for characterizing the corresponding quantum ground states [7, 32, 33, 73, 35]. Moreover, the classical behavior described here could be extended to higher densities, where longer-wavelength mesophases [3] may give rise to even more complex phenomena than those reported in the present study.